Variable Separable: Give the particular and general solution

1. dr/dt = —4rt;

when t =0, r = 0.


2. 2xyy’ =1+y?

when x = 2, y =3


3. xyy’=1+y?

when x =2, y =3.


4. 2ydx = 3xdy

When x =2, y= 1.


5. 2ydx = 3xdy

When x = —2, y= 1.


6. 2ydx = 3xdy;

When x =2, y= —l.


7. y' =x^ (y— x^2);

when x= 0, y =0.


8. xy^2 dx +e^x dy =0;

when x=∞, y=1/2


9. (2a^2 —r^2) dr = r3 sin Ɵ dƟ

When Ɵ =0, r =a.


10.v(dv/dx) = g

when x = xo, v = vo


1
Expert's answer
2021-02-28T07:38:58-0500

Solution.

1)

"\frac{dr}{dt}=-4rt
\newline
\frac{dr}{r}=-4tdt\newline
\ln|r|=-2t^2+C \newline
r=C_1e^{-2t^2}" general solution

When "t=0, r=0," then "C_1=0."

"r=0" particular solution.

2)

"2xyy'=1+y \newline
\frac{y}{y+1}dy=\frac{dx}{2x} \newline
y+1-\ln{|y+1|}=\frac{1}{2}\ln{|x|}+C \newline
y-\ln{|y+1|}=\frac{1}{2}\ln{|x|}+C_1 \newline" general solution

When "x=2, y=3," then

"3-\ln{4}=\ln{\sqrt{2}}+C_1, \newline
C_1=3-\ln{4}-\ln{\sqrt{2}}=3-\ln{\sqrt{32}}."

"y-\ln{|y+1|}=\frac{1}{2}\ln{|x|}+3-\ln{\sqrt{32}}" particular solution.

3)

"xyy'=1+y \newline
\frac{y}{y+1}dy=\frac{dx}{x} \newline
y+1-\ln{|y+1|}=\ln{|x|}+C \newline
y-\ln{|y+1|}=\ln{|x|}+C_1 \newline" general solution

When "x=2, y=3," then

"3-\ln{4}=\ln{2}+C_1, \newline
C_1=3-\ln{4}-\ln{2}=3-\ln{8}."

"y-\ln{|y+1|}=\ln{|x|}+3-\ln{8}" particular solution.

4)

"2ydx=3xdy \newline
2\frac{dx}{x}=3\frac{dy}{y} \newline
2\ln{|x|}=3\ln{|y|}+C \newline
y=C_1x^{\frac{2}{3}}" general solution

When "x=2, y=1," then

"1=C_12^{\frac{2}{3}}, C_1=\frac{1}{\sqrt[3]{4}}=\frac{\sqrt[3]{2}}{2}."

"y=\frac{\sqrt[3]2}{2}x^{\frac{2}{3}}" particular solution.

5)

"2ydx=3xdy \newline
2\frac{dx}{x}=3\frac{dy}{y} \newline
2\ln{|x|}=3\ln{|y|}+C \newline
y=C_1x^{\frac{2}{3}}" general solution

When "x=-2, y=1," then

"1=C_1(-2)^{\frac{2}{3}}, C_1=-\frac{1}{\sqrt[3]{4}}=-\frac{\sqrt[3]{2}}{2}."

"y=-\frac{\sqrt[3]2}{2}x^{\frac{2}{3}}" particular solution.

6)

"2ydx=3xdy \newline
2\frac{dx}{x}=3\frac{dy}{y} \newline
2\ln{|x|}=3\ln{|y|}+C \newline
y=C_1x^{\frac{2}{3}}" general solution

When "x=2, y=-1," then

"-1=C_12^{\frac{2}{3}}, C_1=-\frac{1}{\sqrt[3]{4}}=-\frac{\sqrt[3]{2}}{2}."

"y=-\frac{\sqrt[3]2}{2}x^{\frac{2}{3}}" particular solution.

7)

"y'=x^{y-x^2}"

This is not an equation with separate variables.

Maybe there is an error in the task condition?

8)

"xy^2dx+e^xdy=0 \newline
\frac{xdx}{e^x}=-\frac{dy}{y^2} \newline
-xe^{-x}-e^{-x}=\frac{1}{y}+C \newline
y=-\frac{e^x}{x+1+Ce^x}" general solution

When "x=\infty, y=\frac{1}{2}," then

"C=-{2}."

"y=-\frac{e^x}{x+1-2e^x}" particular solution.

9)

"(2a^2-r^2)dr=r^3\sin{\theta}d\theta \newline
\frac{2a^2-r^2}{r^3}dr=\sin{\theta}d\theta \newline
2a^2\int r^{-3}dr-\int \frac{dr}{r}=\int \sin{\theta}d\theta \newline
-\frac{a^2}{r^2}-\ln{|r|}=-\cos{\theta}+C \newline
\ln{|r|}=\cos{\theta}-\frac{a^2}{r^2}+C_1 \newline" general solution

When "\theta=0, r=a," then

"C_1=\ln{|a|}-1+1=\ln{|a|}."

"\ln{|r|}=\cos{\theta}-\frac{a^2}{r^2}+\ln{|a|}."

10)

"v\frac{dv}{dx}=g \newline
vdv=gdx \newline
\frac{v^2}{2}=gx+C \newline
v^2=2gx+C_1" general solution

When "x=x_0, v=v_0," then

"C_1=v_0^2-2gx_0."

"v^2=2gx+v_0^2-2gx_0, \text{ or} \newline
v=\pm\sqrt{2gx+v_0^2-2gx_0}" particular solution.